OpenStudy (anonymous):

If cos x = c, then sin^2 x/2 equals (1) 1-c / 2 (3) 1-c / c+c (2) 1+c / 2 (4) 1-c / 4

OpenStudy (anonymous):

1

OpenStudy (anonymous):

explanation please

OpenStudy (anonymous):

So this is a given rule for trig but I can try to explain it

OpenStudy (anonymous):

So \[\sin^2x = (1-\cos2x)/2\]

OpenStudy (freckles):

hint: \[\sin^2(x)=1-\cos^2(x)\]

OpenStudy (anonymous):

So given what I just posted when you make it x/2 you just divide all the x's in the thing by 2... So yeah it becomes 1-c/2 if you want more detail from how freckles get's his answer I can help with that too.

OpenStudy (anonymous):

hiiii

OpenStudy (anonymous):

Hello

OpenStudy (anonymous):

I know Terius in real life :) I am the one asking for the help

OpenStudy (anonymous):

Oh okay so do you get my explanation or not?

OpenStudy (anonymous):

Yush and I understand it fully :3 thanks bae

OpenStudy (anonymous):

Uhm no problem =)